OpenStudy (anonymous):

What is the soluton to this problem? 3/2+2b=1

OpenStudy (anonymous):

b = -1/4

OpenStudy (anonymous):

What steps did you use to solve it?

OpenStudy (anonymous):

Subtract 3/2 from both sides of equation Divided both sides of the equation by 2

OpenStudy (anonymous):

Is it possible to multiply 2 on both sides instead of subtracting 3/2?

OpenStudy (anonymous):

Yes, here's how that would look: 2*[3/2+2b=1] 3 + 4b = 2 4b = -1 b = -1/4

OpenStudy (anonymous):

Oh i get it now thanx